OpenStudy (anonymous):

hey ppl ..solve this find l a-b l, lal=2 lbl=3 a.b =4

OpenStudy (y2o2):

\[|a-b| = \sqrt{(a-b)^2} = \sqrt{a^2 -2ab +b^2} = \sqrt{4 -8 +9} = \sqrt{5}\]

OpenStudy (anonymous):

gud 1....

OpenStudy (anonymous):

c d oder ques..

OpenStudy (y2o2):

actually this problem is wrong because a = 2 or -2 b = 3 or 3 so a*b = 6 or -6

OpenStudy (y2o2):

b = 3 or -3 *